Work the integral. 2x + 4x3 dx SVx - 22 TO
Answer #1

a tn Yt1 2 & Sort (No - 2x +4x3) dr = S(60) -2x + 4 x3] du = (6x) dx - 2 (inde +42 m2 de a xt (1+1) (3+1) 2+1 1+1 @jét! 3+1

Please give upvote, if you like the explanation